# Model-guided Fuzzing of Distributed Systems

### v2

@@ 5.3. RedisRaft
RedisRaft [^11] https://github.com/RedisLabs/redisraft powers the popular high-performance Redis distributed key-value store. It compiles into a module that can be loaded onto the main Redis server, which enables different Redis servers to behave as a group and commit client requests in the same order. We instrumented RedisRaft to intercept the exchanged messages [^12] https://anonymous.4open.science/r/instrumented-redisraft-1485/README.md and implement the fuzzer [^13] https://anonymous.4open.science/r/redisraft-fuzzing-E49B/README.md.
We tested RedisRaft running the fuzzer with three processes and five client requests. We ran the test executions with a crash quota of 10 and delivered a maximum of five messages at each step. For each test run, we execute 20K test iterations and report the average results of 20 test runs.
@@
Figure LABEL:fig:redisraft-cov illustrates the average coverage measures for 20K test iterations for each testing strategy. Similar to Etcd, we show that ModelFuzz can obtain better coverage over model states compared to random testing, line coverage guided, and trace coverage-guided fuzzing strategies. On average, ModelFuzz observes 2.58x more states than random exploration, 2.43x more than line coverage, and 2.84x more coverage than trace-guided fuzzing. As with Etcd-Raft, we answer  RQ1 with an observation that model guidance outperforms random exploration and other coverage-guidance strategies in the coverage of explored system states.
We perform the Mann-Whitney U test over the final coverage numbers to mitigate the effect of randomness. We obtain p values of $1.23\mathrm{e}{-4}$ vs random, $8.04\mathrm{e}{-5}$ vs trace and $5.26\mathrm{e}{-4}$ vs line. The U tests conclude that ModelFuzz covers significantly more states than the other testing strategies.
We also analyze the branch coverage of the tests and observe that guiding the test executions using model coverage does not degrade the coverage over traditional code coverage metrics. Table LABEL:tbl:redis-branch-cov reports the mean and standard deviation branch coverage of the different guidance methods. We measure branch coverage of the source code in C using the gcov tool. However, we observe high variance in the data, which we attribute to the coverage measurement tool. As we run multiple copies of the source code during each test iteration, the branch coverage is combined for each copy. However, we found that the gcov tool does not always merge the coverage values of concurrent invocations correctly. Failing to merge coverage information results in the high variance we observe with the branch coverage measures presented in the table.
@@
Our experiments for testing RedisRaft discovered 14 different bugs, two of which are known bugs reported in RedisRaft’s issue tracker, and the remaining 12 are new, previously unknown bugs. The bugs occur in the existence of process crashes and restarts with certain orderings of events, and they manifest as thrown exceptions or assertion violations. We investigated the bugs and reported them in the issue tracker of the RedisRaft open-source repository. [^14] https://github.com/RedisLabs/redisraft/issues (Issue numbers: #643-#649) Table 3 briefly describes the new bugs we discovered.
Table 3. The new bugs found in RedisRaft
@@
8 Process crashes when updating the snapshot index offset
@@
ID ModelFuzz Random Trace Line
1 299(20) 227(20) 368(20) 256(17)
2 10409(15) 13420(13) 8518(11) 7592(10)
3 48(20) 19(20) 32(20) 43(17)
4 10255(17) 12823(18) 11600(18) 10581(14)
5 578(20) 696(20) 945(20) 482(17)
6 8334(3) - - 17784(1)
7 6925(1) 14345(4) - 6512(2)
8 - - 16275(1) -
9 11155(16) 12449(12) 12766(13) 15157(13)
10 11748(2) 6598(3) 18001(1) 9680(2)
11 12031(4) 14041(4) 12158(8) 12261(9)
12 5709(1) 11832(2) 16097(1) -
13 6563(1) - - -
14 862(1) - - -
ID Random Trace Line
1 0.364 0.504 0.462
2 0.641 0.412 0.373
3 0.373 0.433 0.443
4 0.559 0.569 0.563
5 0.545 0.455 0.522
6 1.000 1.000 1.000
7 0.750 1.000 0.500
8 N/A 0.000 N/A
9 0.547 0.567 0.611
10 0.167 1.000 0.250
11 0.563 0.531 0.500
12 1.000 1.000 1.000
13 1.000 1.000 1.000
14 1.000 1.000 1.000
@@
Figure 6. (a) The number of RedisRaft tests for the first occurrences of the bugs using different guidance strategies. (b) Pairwise $\hat{A}_{12}$ statistic results against ModelFuzz.
@@
Among all 14 bugs, ModelFuzz found more bugs than other guidance approaches. Specifically, ModelFuzz found 13 bugs, while random exploration, trace-guided, and line-guided found only 10 of them. Furthermore, ModelFuzz is faster in reproducing 7 of the bugs compared to other approaches, where Random exploration and Line-guided each find three, and trace-guided finds one bug faster. For each bug and guidance method, the table in Figure LABEL:tbl:redis-bugs-alt lists the number of runs that find the bug (in parenthesis) and the average first occurrence of the bug in the successful runs. For each bug, we highlight the lowest average first occurrence in bold. Among the bugs, three of them are found only using ModelFuzz. We classify a bug as rare if it was found in at most five trials. ModelFuzz finds four of the rare bugs faster than the other approaches.
@@
For each of the 14 bugs, we calculate Vargha and Delaney’s $\hat{A}_{12}$ statistics to analyze the pair-wise statistical significance of the results. Table LABEL:tbl:redis-bug-stats reports the $\hat{A}_{12}$ statistics against ModelFuzz for each bug, highlighting the results with statistical significance in bold. The analysis shows that ModelFuzz detects the bugs ${\{6,12,13,14\}}$ statistically significantly faster than all other guidance approaches, ${\{2,7,9,10\}}$ statistically significantly faster than some of the approaches, and comparably faster for the remainder. While this indicates its ability to detect bugs faster, the $\hat{A}_{12}$ results do not draw clear conclusions on the statistical significance.
Moreover, we test for statistical significance across all bugs in the mean iterations to find the bugs. Since we compare more than two approaches and the number of iterations for all bugs are normal and homoscedastic, we use repeated measures ANOVA (girden1992anova, ) as omnibus tests and post-hoc Tukey HSD (tukey1949comparing, ) for assessing overall significance. We reject the null hypothesis ($p=0.008$) of the repeated measure ANOVA, as there is a statistically significant difference between the mean values of the approaches. Overall, the tests show that ModelFuzz statistically significantly detects bugs faster.
Figure 7(a) reports the mean, standard deviation, and confidence intervals for all approaches, and Figure 7(b) illustrates the Tukey HSD results. We can observe that the mean iteration to detect a bug for ModelFuzz lies outside of the other approaches. This suggests that the differences between ModelFuzz and all of the remaining approaches are statistically significant, whereas there is no significant difference between random exploration, line coverage, and trace coverage guidance. Note that we perform ANOVA and post-hoc Tukey HSD tests only for RedisRaft, as we analyze numerous bugs with which we could form a sample set.
Method Mean $\pm$ std $CI_{l}$ $CI_{u}$
ModelFuzz 7494 $\pm$ 5742 3922 11066
Line 11454 $\pm$ 7675 7881 15025
Random 11890 $\pm$ 7377 8318 15462
Trace 12626 $\pm$ 7520 9054 16198
Figure 7. (a) Mean, standard deviation, lower confidence ($CI_{l}$), and upper confidence interval values ($CI_{u}$) of the first iteration to find a bug. (b) The Tukey HSD test plot.
@@ 5.4. Comparison to related work
Close to our approach, recent work Mallory (DBLP:journals/corr/abs-2305-02601, ) uses greybox fuzzing to test distributed system implementations. Mallory utilizes reinforcement learning to select the next events using a timeline abstraction of the system states, which is similar to the trace coverage notion. However, the set of traces is typically too large for effective guidance. Mallory overcomes this problem by allowing the user to manually annotate the system’s code to specify the interesting parts.
@@
Our evaluation does not provide an empirical comparison to Mallory for all the systems we tested since the application of Mallory requires manual inspection and annotation of the source code under test. However, we share the RedisRaft case study with Mallory, and we compare our empirical results for testing RedisRaft with that of Mallory, as presented in their work (DBLP:journals/corr/abs-2305-02601, ).
@@
Our tests replicate 2 of the previously known bugs found by both Jepsen and Mallory, and we find 6 new bugs in RedisRaft that were previously not found (described in Table 3). The existence of bugs that could only be detected by Mallory or only by ModelFuzz suggests the complementary bug-finding capabilities of the two methods. ModelFuzz may miss bugs in parts of the implementation (e.g., rolling back a snapshot) that are not captured by the system’s abstract model used. On the other hand, it can uncover bugs that occur in the parts of the system code that are not user-annotated but captured in the model.
@@ 6. Limitations
Model-guided fuzzing requires the model to be close to the implementation so that the model provides effective feedback for test generation. The abstraction level of the model heavily affects the performance of model guidance. A high-level model abstracting away critical aspects will be ineffective at guiding the tests toward exploring these aspects in the system implementation. For example, while the TLA+ model of Raft protocol (url/raft-tla, ) provided by the protocol’s authors models process interactions, it does not capture process crash and restart actions. Therefore, using this model to guide test generation will be ineffective at directing the test executions with process crashes and restarts that potentially lead to previously unseen executions. Conversely, a model that is too fine-grained and captures all implementation details is also not helpful. For example, a model that implements HeartBeat messages or captures too many details of process states distinguishes equivalent high-level behaviors of the system. It can guide test generation toward exploring different executions that produce equivalent system behavior.
@@ 7. Related Work
Coverage-guided fuzzing  (DBLP:conf/ccs/BohmePR16, ; DBLP:conf/kbse/LemieuxS18, ; zeller2019fuzzing, ; DBLP:journals/tse/ManesHHCESW21, ; DBLP:conf/uss/BaBMR22, ; Heuse_AFL_2022, ) has been extensively studied for test input generation for sequential programs. Recent techniques target the generation of different types of program inputs (DBLP:conf/sigsoft/GodefroidHP20, ; DBLP:journals/corr/abs-2109-11277, ; DBLP:conf/sigsoft/SteinhofelZ22, ) and improve fuzzer performance (DBLP:conf/sigsoft/BohmeMC20, ; DBLP:conf/issta/ShouTS23, ). Extensions of AFL  (url/afl, ) such as AFLNet (DBLP:conf/icst/PhamBR20, ), StateAFL (DBLP:journals/ese/Natella22, ) test communication protocols by mutating structured message inputs guided by the states explored. Different from test input fuzzing, we utilize the fuzzing approach to generate event schedules.
@@
Testing distributed systems has been the focus of a wide range of research such as systematic testing (DBLP:conf/nsdi/YangCWXLLYLZZ09, ; DBLP:conf/spin/SimsaBG11, ; DBLP:conf/sigsoft/DesaiQS15, ; DBLP:conf/fast/DeligiannisMTCD16, ; DBLP:conf/osdi/LeesatapornwongsaHJLG14, ; DBLP:conf/asplos/Leesatapornwongsa16, ; DBLP:journals/pacmpl/PickDG23, ), random fault injection (url/jepsen, ) or scheduling (url/namazu, ), and designing methods for carefully sampling event schedules and faults (DBLP:conf/cav/ChistikovMN16, ; DBLP:journals/pacmpl/OzkanMNBW18, ; DBLP:journals/pacmpl/OzkanMO19, ; DBLP:conf/asplos/YuanY20, ; DBLP:journals/pacmpl/DragoiEOMN20, ; DBLP:journals/pacmpl/WinterBGGO23, ). While some methods guide fault injection using lineage (DBLP:conf/sigmod/AlvaroRH15, ), runtime state (DBLP:conf/kbse/ChenDWQ20, ), or system’s meta-variables (DBLP:conf/sosp/LuL00TYY19, ), most techniques do not exploit information observed in the test executions and generate tests independently from each other.
Fuzzing concurrent and distributed systems. Fuzzing methods for multithreaded concurrency guide the tests by monitoring races and synchronization events (DBLP:conf/pldi/Sen08, ; DBLP:conf/icse/WangSG11, ; DBLP:conf/oopsla/YuNPP12, ), execution states caused by thread interleavings (DBLP:conf/uss/ChenGXSZLWL20, ), coverage of concurrent call pairs (DBLP:conf/ndss/JiangBL022, ), and using the reads-from relation between the memory accesses (conf/asplos/WZDMR24, ). These methods are designed for multithreaded programs, and they do not target distributed concurrency. Recent testing techniques learn from the explored executions and adapt reinforcement learning or fuzzing approaches for the generation of new tests. QL (DBLP:journals/pacmpl/MukherjeeDBL20, ) uses reinforcement learning to guide the exploration to unexplored parts of the execution space. Evolutionary search-based testing (DBLP:conf/icse/MeertenOP23, ) directs the exploration toward a fitness function. Existing work for fuzzing distributed systems uses code coverage and message traces as program feedback. CrashFuzz (DBLP:conf/icse/GaoDWFWZH23, ) and FaultFuzz(DBLP:conf/icse/FengP0WD0LL24, ) fault-injection tools inject process crashes using code coverage information as program feedback. Mallory (DBLP:journals/corr/abs-2305-02601, ) builds on Jepsen (url/jepsen, ) and guides test generation using programmer annotations that mark interesting parts of the code and an event timeline abstraction (close to traces) as the feedback information.
@@
Model-based testing uses formal models (e.g., TLA+ specifications) to exhaustively enumerate system executions and enforce them on the implementation. Its applications include testing APIs (DBLP:conf/hvc/ArthoBHPSTY13, ), fragments of HTTP protocol (DBLP:conf/issta/LiPZ21, ) and specific systems (DBLP:journals/pvldb/SchvimerDH20, ). Protocol fuzzers DTLS-Fuzzer (DBLP:conf/icst/Fiterau-Brostean22, ) and EDHOC-Fuzzer (DBLP:conf/issta/SagonasT23, ) use model learning to generate a state machine model of the protocol implementations which can be used for model-based testing.
@@
Recent works Mocket (DBLP:conf/eurosys/WangDGWW023, ) and MET (DBLP:journals/smr/ZhangHWM24, ) adopt model-based testing to test distributed systems and CRDTs. Mocket uses the paths in the model’s state space graph as test cases, and it enforces the system under test to run the actions generated on the system’s model on the corresponding states and actions in the implementation. For this, it synchronizes the executions of the system and the model at each step, which requires heavy annotation on the system’s source code. Our evaluation does not empirically compare to Mocket since we do not have the annotations to map the source codes of the systems under test to the TLA+ specifications.
@@
ModelFuzz conceptually differs from model-based testing, as it performs an unconstrained exploration of the implementation. Model-based testing generates test cases using model paths, and hence, it does not cover parts of the implementation that are abstracted away in the model. In contrast, model-guided fuzzing explores the executions of the implementation, including those not captured by the model.
@@
Model-guided fuzzing stands out from the state-of-the-art in testing distributed systems as it steers the exploration towards more coverage of system behaviors without the need for the programmer’s comprehension and annotations on the source code. Instead, it utilizes the system’s abstract model, which already summarizes the essential information about system behavior, to collect information about the explored system behaviors and generate new tests.

### v3

@@ 5.4. RedisRaft
RedisRaft [^16] https://github.com/RedisLabs/redisraft powers the popular high-performance Redis distributed key-value store. It compiles into a module that can be loaded onto the main Redis server, which enables different Redis servers to behave as a group and commit client requests in the same order. We instrumented RedisRaft to intercept the exchanged messages [^17] https://github.com/egeberkaygulcan/redisraft-fuzzing and implement the fuzzer [^18] https://github.com/zeu5/redisraft-fuzzing. Overall, the Raft module consists of 30k lines of C code. Our instrumentation adds an additional 1.5k lines of C and Go code, where the Go code implements the fuzzer loop. Unlike with Etcd, the instrumentation of RedisRaft requires running all the nodes in different OS processes, along with a network server to capture and deliver messages. Therefore, starting, stopping, and restarting nodes is a time-consuming process.
We tested RedisRaft by running the fuzzer with three processes and five client requests. We ran the test executions with a crash quota of 10 and delivered a maximum of five messages at each step. For each test run, we executed the test for 12 hours, each iteration running for a fixed duration of 3s, and reported the average results of 21 test runs.
@@
Figure LABEL:fig:redisraft-cov illustrates the average coverage measures for 12-hour test runs for each testing strategy. We show that ModelFuzz can obtain better coverage over model states compared to random testing, line coverage-guided, and trace coverage-guided fuzzing strategies. The coverage of the reinforcement learning-based approach BonusMaxRL is comparable to ModelFuzz.
On average, ModelFuzz observes 2.58x more states than random exploration, 2.43x more than line coverage, 2.84x more coverage than trace-guided fuzzing, and 1.06x more than BonusMaxRL. We answer  RQ1 with an observation that model guidance outperforms random exploration and other coverage-guidance strategies in the coverage of explored system states and achieves comparable coverage to the BonusMaxRL strategy.
We perform the Mann-Whitney U test over the final coverage numbers to mitigate the effect of randomness. We obtain p values of $3.12\mathrm{e}{-8}$ vs random, trace, and line coverage guidance strategies. The U tests show that ModelFuzz covers significantly more states than the other traditional guidance strategies, barring the reinforcement learning guided approach.
We also analyze the branch coverage of the tests and observe that guiding the test executions using model coverage does not degrade the coverage over traditional code coverage metrics. Table LABEL:tbl:redis-branch-cov reports the mean and standard deviation branch coverage of the different guidance methods. We measure the branch coverage of the source code in C using the gcov tool. However, we observe a high variance in the data, which we attribute to the coverage measurement tool. As we run multiple copies of the source code during each test iteration, the branch coverage is combined for each copy. However, we found that the gcov tool does not always merge the coverage values of concurrent invocations correctly. Failing to merge coverage information results in the high variance we observe with the branch coverage measures presented in the table.
Unlike for Etcd-raft, ModelFuzz is able to cover more states when compared to other approaches. The difference lies in the nature of the instrumentation. The performance overhead (in time) of communicating with TLC to obtain coverage information is insignificant compared to the time required to restart all the nodes. Therefore, all approaches cover a similar number of iterations, and we are able to achieve better coverage.
@@
Our experiments for testing RedisRaft discovered 13 different bugs, two of which are known bugs reported in RedisRaft’s issue tracker, and the remaining 11 are new, previously unknown bugs. The bugs occur in the existence of process crashes and restarts with certain orderings of events, and they manifest as thrown exceptions or assertion violations. We investigated the bugs and reported them in the issue tracker of the RedisRaft open-source repository. [^19] https://github.com/RedisLabs/redisraft/issues (Issue numbers: #643-#649) Table 3 briefly describes the new bugs we discovered.
@@
Table 3. The new bugs found in RedisRaft
ID Random Trace Line RL ModelFuzz
1 0.12(21) 0.18(21) 0.15(21) - 0.14(21)
2 3.72(19) 4.26(20) 4.06(15) 0.02(1) 4.13(21)
3 0.01(21) 0.01(21) 0.01(21) 0.02(21) 0.02(21)
4 4.84(16) 5.00(13) 3.45(10) 3.15(3) 3.63(12)
5 1.78(21) 1.35(21) 1.41(21) 5.91(8) 1.75(21)
6 - - - - 9.8(1)
7 - 8.76(2) 5.21(1) - -
8 1.94(21) 1.82(21) 0.93(21) 2.68(19) 1.14(21)
9 5.76(6) 5.05(4) 2.78(5) - 4.71(5)
10 - - - 0.17(1) 3.89(1)
11 4.67(1) 9.60(1) - - 11.41(1)
12 - - - - 3.93(1)
13 - - - - 0.51(1)
ID Random Trace Line RL
1 0.399 0.467 0.494 1.000
2 0.491 0.540 0.521 0.000
3 0.434 0.506 0.471 0.585
4 0.529 0.604 0.524 0.462
5 0.480 0.435 0.465 0.892
6 1.000 1.000 1.000 1.000
7 - - - -
8 0.628 0.637 0.439 0.667
9 0.583 0.5 1.000 1.000
10 1.000 1.000 1.000 0.000
11 0.000 0.000 1.000 1.000
12 1.000 1.000 1.000 1.000
13 1.000 1.000 1.000 1.000
@@
Figure 7. (a) The number of RedisRaft tests for the first occurrences of the bugs (in hours) using different guidance strategies. (b) Pairwise $\hat{A}_{12}$ statistic results against ModelFuzz.
@@
Among all 13 bugs, ModelFuzz found more bugs than other guidance approaches. Specifically, ModelFuzz found 12 bugs, while random exploration found 8, trace-guided found 9, line-guided found 8, and BonusMaxRL found only 6 of them. For each bug and guidance method, the table in Figure LABEL:tbl:redis-bugs-alt lists the number of runs that find the bug (in parenthesis) and the average first occurrence of the bug (in hours) in the successful runs. For each bug, we highlight the lowest average first occurrence in bold. Among the bugs, one of them is found only using ModelFuzz. Overall, ModelFuzz is more consistent in reproducing the bugs, finding them more frequently than other approaches.
@@
For each of the bugs, we calculate Vargha and Delaney’s $\hat{A}_{12}$ statistics to analyze the pairwise statistical significance of the results. Table LABEL:tbl:redis-bug-stats reports the $\hat{A}_{12}$ statistics against ModelFuzz for each bug, highlighting the results with statistical significance in bold. The analysis shows that ModelFuzz detects the bug $\{6\}$} statistically significantly faster than all other guidance approaches, $\{8,9,10,11\}$} statistically significantly faster than some of the approaches, and comparably faster for the remainder. While this indicates its ability to detect bugs faster, the $\hat{A}_{12}$ results do not draw clear conclusions on the statistical significance.
What is significant is the number of bugs and the occurrences of the bugs found by ModelFuzz compared to other approaches. Specifically, while RL based approach BonusMaxRL achieves comparable coverage over the states, the number of bugs found by BonusMaxRL is significantly lesser than ModelFuzz or other traditional guidance measures. The targeted coverage guided fuzzing approach is able to lead the exploration to corner cases leading to better bug finding ability.
@@ 5.5. Comparison to Related Work
5.5.1. Comparison to Mallory
A closely related work to our approach is Mallory (Meng et al., 2023), a greybox fuzzer for distributed systems that employs reinforcement learning to guide test generation.
While both Mallory and ModelFuzz learn from test executions to guide test generation, they adopt different approaches to learning. Mallory runs a long test execution (with a reported duration of 24 hours in their paper) and learns from the exploration within that test execution. After each action taken during the execution, it updates its reinforcement learning policy and makes decisions about subsequent actions based on that information. On the other hand, ModelFuzz runs short test executions (each with a duration of a few seconds) and learns from the exploration across executions. After each test, it analyzes the model coverage achieved and generates the next test cases by mutating the previous executions that have led to coverage of new states.
@@ 5.5.1. Comparison to Mallory
The two approaches also differ in the coverage information they use to learn from the executions. Mallory uses a timeline abstraction of the system states, which is similar to the trace coverage notion. However, the set of traces is typically too large for effective guidance. Mallory overcomes this problem by allowing the user to manually annotate the system’s code to specify the interesting parts and use this information in the learning process. On the other hand, ModelFuzz uses the abstract state coverage based on the formal model states (in our experiments, TLA+) of a system.
@@
Our evaluation does not provide an empirical comparison to Mallory across all the benchmark systems we tested, since the application of Mallory requires manual inspection and annotation of the source code under test. Therefore, we focus our comparison on the shared benchmark, RedisRaft. Specifically, we compare the abstract state coverage and bug detection of the two techniques.
We ran Mallory using the default parameters in its artifact repository, updating its configuration files for the Docker setup and adjusting for differences in our test platforms. Because Mallory’s standard logs lack the detailed event and state information needed to measure abstract state coverage, we extended its instrumentation to RedisRaft to log this information. A limitation in measuring the abstract state coverage of Mallory tests is due to the length of their test executions. A single test execution of Mallory processes thousands of user requests, as opposed to five client requests in ModelFuzz tests, making it infeasible to run TLC simulations on the event traces. We mitigate this by collecting abstract state information during the test execution, at the same points where Mallory collects its event information. Since Mallory collects the process logs separately, rather than compiling them into a combined serial log of system events like ModelFuzz does, we merged the process logs of Mallory executions, ordering them by their timestamps. This allowed us to reconstruct the sequence of system states explored during execution. We read the logs into the same definition of model state in Mallory as we used for ModelFuzz.
Figure 8. Comparison of abstract state coverage of Mallory and ModelFuzz for testing Redis for 8-hour test runs with the state abstraction of term numbers (a) including the set of active processes (b) excluding the set of active processes in the abstract system state.
State coverage.
We ran Mallory and ModelFuzz for an 8-hour duration of testing, repeated the tests 20 times, and reported their averages. To ensure a fair comparison in the number of requests processed by Mallory and ModelFuzz, we normalized the logs so that the processes process the same maximum number of user requests simultaneously in the executions of both methods.
The coverage results showed a linear increase in Mallory’s state coverage over the duration of fuzzing. Upon analyzing the covered states, we observed that the linear growth in coverage is introduced by the increasing term numbers during the protocol execution. The term number is a monotonically increasing counter that tracks the leader election processes in the protocol execution. During the 8-hour test execution of Mallory, the term number increases with the test execution time, while it remains relatively constant, within a few seconds of ModelFuzz tests. To eliminate the effect of the term counter in counting distinct states, we abstracted the term numbers in the state information by considering two states that differ only in the term number as equivalent.
Figure LABEL:fig:modelfuzz-mallory-with-term-abst presents the abstract state coverage of Mallory and ModelFuzz with the term number abstraction. The results show that ModelFuzz and Mallory perform comparably, with ModelFuzz achieving higher coverage. Figure LABEL:fig:modelfuzz-mallory-ignored-active presents the state coverage with an additional abstraction of the information of the set of currently active processes in the system state. As given in Section 4.2.1, we extend Raft’s TLA+ model to track the information of crash and restart of the processes and keep that information as part of the abstract system state. ModelFuzz utilizes this information to guide the new tests through the controlled injection of process crashes. However, the exact points of process crashes and restarts are not available in the Mallory logs and, therefore, are not read into the abstract states. To present a fair comparison based on the same set of collected data, we additionally present the state coverage results without this information in Figure LABEL:fig:modelfuzz-mallory-ignored-active. The results show comparable performance for Mallory and ModelFuzz.
Bug finding.
@@
ModelFuzz tests replicate 2 of the previously known bugs found by both Jepsen and Mallory, and we find 6 new bugs in RedisRaft that were previously not found (described in Table 3). The existence of bugs that could only be detected by Mallory or only by ModelFuzz suggests the complementary bug-finding capabilities of the two methods. ModelFuzz may miss bugs in parts of the implementation (e.g., rolling back a snapshot) that are not captured by the system’s abstract model used. On the other hand, it can uncover bugs that occur in the parts of the system code that are not user-annotated but captured in the model.
Overall, Mallory and ModelFuzz show comparable performance in abstract state coverage and bug detection, and they can be used complementarily to each other. The performance of Mallory is influenced by user annotations that identify significant states, while ModelFuzz relies on the expressiveness of the abstract model used. Depending on the user annotations for Mallory and the abstract model used by ModelFuzz, they can be more effective at covering and detecting bugs in the complementary parts of the execution space.
5.5.2. Comparison to BonusMaxRL
We empirically compare our approach against BonusMaxRL (Borgarelli et al., 2024), which utilizes reinforcement learning techniques to explore the state space of a distributed system. Specifically, BonusMaxRL provides a novel reward mechanism that extends prior work and has been proven to be successful in finding bugs in both RedisRaft and Etcd-raft benchmarks. We present empirical comparison of the coverage for all our benchmarks. For the Etcd benchmark, BonusMaxRL outperforms ModelFuzz while for the RedisRaft benchmark, BonusMaxRL achieves comparable coverage to ModelFuzz. However, we observe that BonusMaxRL fails to find all the bugs in both benchmarks.
5.6. Summary of the Evaluation Results
In summary, we answer RQ1, showing that ModelFuzz outperforms pure random testing and code-coverage and trace-guided fuzzing strategies in the coverage of model states, yet producing coverage comparable to or worse than BonusMaxRL. We answered RQ2 by comparing the number of test executions until the first occurrence of a bug. We showed that ModelFuzz discovers bugs faster than the other strategies on average.
Moreover, the intuition of using model state coverage to guide the fuzzer has shown effective, discovering 1 new bug for Etcd-raft and 11 new bugs for RedisRaft, along with replicating 2 previously known bugs. Note that 2 of the new bugs in RedisRaft discovered by ModelFuzz could not be detected by other approaches in the given test budget, showing that our approach is more robust in finding bugs earlier than other guidance approaches.
Finally, our evaluation on several systems under test that implement different distributed protocols demonstrates the applicability of ModelFuzz across multiple systems and the generalizability of the findings.
@@ 6. Limitations
Model-guided fuzzing requires the model to be close to the implementation so that the model provides effective feedback for test generation. The abstraction level of the model heavily affects the performance of model guidance. A high-level model abstracting away critical aspects will be ineffective at guiding the tests toward exploring these aspects in the system implementation. For example, while the TLA+ model of Raft protocol (Github., [n. d.]) provided by the protocol’s authors models process interactions, it does not capture process crash and restart actions. Therefore, using this model to guide test generation will be ineffective at directing the test executions with process crashes and restarts that potentially lead to previously unseen executions. Conversely, a model that is too fine-grained and captures all implementation details is also not helpful. For example, a model that implements HeartBeat messages or captures too many details of process states distinguishes equivalent high-level behaviors of the system. It can guide test generation toward exploring different executions that produce equivalent system behavior.
@@ 7. Related Work
Fuzzing (Böhme et al., 2016; Lemieux and Sen, 2018; Zeller et al., 2019; Manès et al., 2021; Ba et al., 2022; Heuse et al., 2022) has been extensively studied for test input generation for sequential programs. Rather than generating test inputs independently at random, coverage-guided fuzzing techniques incorporate lightweight program instrumentation to collect some feedback information from the test executions and use that information to guide the generation of new tests. Recent techniques target the generation of different types of program inputs (Godefroid et al., 2020; Dutra et al., 2024; Steinhöfel and Zeller, 2022) and improve fuzzer performance (Böhme et al., 2020; Shou et al., 2023). Extensions of AFL  (Zalewski, [n. d.]) such as AFLNet (Pham et al., 2020), StateAFL (Natella, 2022) test communication protocols by mutating structured message inputs guided by the states explored. Different from test input fuzzing, our work utilizes the fuzzing approach to generate event schedules for testing distributed system executions.
Testing distributed systems
@@
Distributed systems have been the focus of a wide range of research such as systematic testing and model checking (Yang et al., 2009; Simsa et al., 2011; Desai et al., 2015; Deligiannis et al., 2016; Leesatapornwongsa et al., 2014, 2016; Pick et al., 2023; Enea et al., 2024), random fault injection (Kingsbury., 2022) or scheduling (GitHub, [n. d.]), and designing methods for carefully sampling event schedules and faults (Chistikov et al., 2016; Kulahcioglu Ozkan et al., 2018, 2019; Yuan and Yang, 2020; Dragoi et al., 2020; Winter et al., 2023). While some methods guide fault injection using lineage (Alvaro et al., 2015), runtime state (Chen et al., 2020a), or system’s meta-variables (Lu et al., 2019), most techniques do not exploit information observed in the test executions and generate tests independently from each other.
Fuzzing concurrent and distributed systems.
Fuzzing methods for multithreaded concurrency guide the tests by monitoring races and synchronization events (Sen, 2008; Wang et al., 2011; Yu et al., 2012), execution states caused by thread interleavings (Chen et al., 2020b), coverage of concurrent call pairs (Jiang et al., 2022), and using the reads-from relation between the memory accesses (Wolff et al., [n. d.]). These methods are designed for multithreaded programs, and they do not target distributed concurrency.
Recent testing techniques for distributed systems learn from the explored executions and adapt reinforcement learning or fuzzing approaches to generate new tests. QL (Mukherjee et al., 2020) uses reinforcement learning to guide the exploration to unexplored parts of the execution space. Recent work (Borgarelli et al., 2024) improves the reward augmentation using theoretical results in reinforcement learning and programmer-provided information for the interesting parts of the search space. Evolutionary search-based testing (van Meerten et al., 2023) directs the exploration toward a fitness function.
Existing work for fuzzing distributed systems uses code coverage and message traces as program feedback. CrashFuzz (Gao et al., 2023) and FaultFuzz(Feng et al., 2024) fault-injection tools inject process crashes using code coverage information as program feedback. Mallory (Meng et al., 2023) builds on Jepsen (Kingsbury., 2022) and guides test generation using programmer annotations that mark interesting parts of the code and an event timeline abstraction (close to traces) as the feedback information. Different from using user annotations, code coverage, or trace coverage for guidance, ModelFuzz uses guidance from the abstract system model.
Model-based testing
@@
Model-based testing uses formal models (e.g., TLA+ specifications) to exhaustively enumerate system executions and enforce them on the implementation. Its applications include testing APIs (Artho et al., 2013), fragments of HTTP protocol (Li et al., 2021) and specific systems (Schvimer et al., 2020). Protocol fuzzers DTLS-Fuzzer (Fiterau-Brostean et al., 2022) and EDHOC-Fuzzer (Sagonas and Typaldos, 2023) use model learning to generate a state machine model of the protocol implementations which can be used for model-based testing.
@@
Mocket (Wang et al., 2023) and MET (Zhang et al., 2024) adopt model-based testing to test distributed systems and CRDTs. Mocket uses the paths in the model’s state space graph as test cases, and it enforces the system under test to run the actions generated on the system’s model on the corresponding states and actions in the implementation. For this, it synchronizes the executions of the system and the model at each step. The synchronization of executions requires heavy annotation on the system’s source code to mark the program variables and messages associated with the system’s model variables.
@@
Our work conceptually differs from model-based testing, as it performs an unconstrained exploration of the implementation. Model-based testing generates test cases using model paths, and hence, it does not cover parts of the implementation that are abstracted away in the model. In contrast, model-guided fuzzing explores the executions of the implementation, including those not captured by the model.
Recent work (Cirstea et al., 2024) uses TLA+ models to validate execution traces of distributed systems. It uses programmer instrumentation on the system implementation to identify the points of state variable updates and transitions and validates the collected execution traces on the TLA+ model to detect the conformance or discrepancies between the system specification and its implementations. While this work focuses on verifying execution traces, its future work points to using TLA+ models for guiding test executions towards interesting parts of the search space, as we do with ModelFuzz.
Model-guided fuzzing
@@
ModelFuzz stands out from the state-of-the-art in testing distributed systems as it steers the exploration towards more coverage of system behaviors without the need for the programmer’s comprehension and annotations on the source code (e.g., to identify and mark interesting parts of the source code or using heavy instrumentation to synchronize its execution with a formal model). Instead, it utilizes the system’s abstract model, which already captures the essential information about system behavior, to collect information about the explored system behaviors and generate new tests.